A body of mass 0.6kg is thrown vertically upward from ground with a speed of 20ms-1. Calculate its potential energy at the maximum height


1
Expert's answer
2022-03-24T16:18:58-0400

The potential energy at the maximum height is equal to the initial kinetic energy

"E_p=\frac{mv_0^2}{2}=\frac{0.6*20^2}{2}\\
=120\:\rm J"